A local garage uses two types of inputs, capital and labour, to service used cars. Their production function is described by Q = 8K^0.8L^0.2, where Q is the number of cars serviced, K is the amount of capital, equipment and tools used in the process and L denotes the input of human labour. The costs of inputs are £16 per unit of labour and £1 per unit of capital. The garage has received an order from a taxi company to service a fleet of 48 cars. Consider a constrained cost minimisation problem and answer the following questions.

a. What is the cost minimising quantity of capital the garage should use?
K* =
b. What is the cost minimising quantity of labour the garage should use?
L* =
c. What is the lowest cost at which 48 cars can be serviced?
1
Expert's answer
2019-07-02T08:52:26-0400

a. The cost minimising quantity of capital the garage should use is at MPK/r = MPL/w, so:

"MPK = Q'(K) = 6.4K^{-0.2}L^{0.2},"

"MPL = Q'(L) = 1.6K^{0.8}L^{-0.8},"

"6.4K^{-0.2}L^{0.2}\/1 = 1.6K^{0.8} L^{-0.8}\/16,"

K = 64L,

"8\u00d7(64L)^{0.8}L^{0.2}= 48,"

L = 0.215,

K = 0.215×64 = 13.78.

K* = 13.78.

b. The cost minimising quantity of labour the garage should use is:

L* = 0.215.

c. The lowest cost at which 48 cars can be serviced is: 0.215×16 + 13.78×1 = 17.22.